# On the Challenges and Opportunities in Generative AI

### v3

@@ 1 Introduction
The past few years have demonstrated the immense potential of large-scale generative models to create powerful AI tools capable of impacting society profoundly. Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Rae et al., 2021) and their dialogue agents, such as ChatGPT (OpenAI, 2023) and Llama 3 (Grattafiori et al., 2024). have enabled the development of highly effective text generation systems that produce coherent, contextually relevant, and user-tailored outputs across a wide range of use cases. Similarly, advancements in diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) have led to groundbreaking advancements in image synthesis tasks, such as large-scale text-to-image generation (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Esser et al., 2024). These successes show that highly effective AI systems can be built using a relatively straightforward recipe: combining simple generative modeling paradigms (Larochelle & Murray, 2011; Sohl-Dickstein et al., 2015) with successful network architectures (Vaswani et al., 2017; Dosovitskiy et al., 2020; Ronneberger et al., 2015), training on large-scale datasets, and the incorporation of preferences via human feedback (Ouyang et al., 2022; Ziegler et al., 2019). The impact of generative AI has not been limited to text and image generation applications. It has fueled accelerated progress across a variety of research fields and practical applications, spanning from biology (Jumper et al., 2021) to weather forecasting (Ravuri et al., 2021), from code generation (Chen et al., 2021b; Li et al., 2022b) to video creation (Yang et al., 2023a; Ho et al., 2022b; Singer et al., 2022; Brooks et al., 2024), audio synthesis (Borsos et al., 2023; Liu et al., 2023a), and even artistic and musical composition (Huang et al., 2023b).
@@
With the current advancements and excitement surrounding generative AI, a question naturally arises: Are we on the brink of an AI utopia? Are we close to defining a perfect generative model—capable of modeling (i) any joint and conditional distribution over real-world data and (ii) its underlying causal structure—that could theoretically solve every conceivable AI task, from discriminative challenges to reinforcement learning and beyond? We argue in this paper that the answer is a resounding no; rather, the realization of such a model, one that would fundamentally transform the field of AI, is still a distant vision, hampered by substantial theoretical, practical, and ethical challenges.
@@
Amidst the excitement and anticipation surrounding this new wave of Deep Generative Models (DGMs), [^1] In this paper, we refer to Generative AI as a collection of large-scale DGMs and use the term DGM henceforth. it is easy to overlook the new set of challenges they introduce. Unlike many of the traditional machine learning models, DGMs generate outputs in very high-dimensional spaces, which introduces several technical complexities. These include significantly increased computational demands, a need for larger datasets to accurately capture the underlying data distribution, and challenges in effectively evaluating and interpreting the generated outputs. And while significant progress has been made in improving interpretability and computational efficiency for traditional models (Marcinkevics & Vogt, 2020), these existing methods are frequently ill-suited for DGMs, at least in part because of the complex and high-dimensional nature of their outputs. Consequently, there is a pressing need for the development of a new set of techniques and tools tailored to these models, particularly to enable efficient inference, interpretability and quantization. These challenges lead us to conclude that scaling up current paradigms is not in isolation the ultimate path towards a perfect generative model. While increasing model size and training data can enhance performance on benchmarks, it does little to address the fundamental shortcomings of DGMs, such as their inefficiency, lack of inclusivity, limited transparency, and barriers to usability—particularly in high-stakes domains where reliability and fairness are paramount.
@@
This paper emerged as a result of the Dagstuhl Seminar on Challenges and Perspectives in Deep Generative Modeling [^2] https://www.dagstuhl.de/23072 held in Spring 2023. By outlining a comprehensive roadmap of the current state and open challenges of generative AI, we hope to empower researchers and practitioners alike, fostering the development of generative AI models that are not only more robust and reliable but also accessible to a wider audience.
@@ 2.1 Generalization and Robustness
To ensure reliability across various domains, DGMs must generalize effectively to shifts in data, also known as out-of-distribution (OOD) robustness, and be resilient to minor variations in the input, i.e., adversarial robustness. Without proper generalization, generative models may produce unrealistic or biased outputs, [^3] Here, we use the terminology biased outputs to refer to systematic deviations in model outputs caused by imbalances or inaccuracies in the training data and/or modeling process. These outputs then do not accurately reflect the true underlying data distribution or are skewed in ways that perpetuate inaccuracies, stereotypes, or unfair conceptions about certain outcomes. limiting their practical utility and trustworthiness in real-world applications.
@@
While large-scale generative models show some promise in achieving OOD robustness (Wang et al., 2023a), these models still face challenges in accurately capturing rare events or responding to adversarial inputs (Zhu et al., 2024), a difficulty that lies in effectively modeling the tail of information (Kandpal et al., 2023), i.e., the information that appears rarely or only once in the dataset used to (pre)train the model. This limitation indicates a gap in their ability to fully represent the vast and diverse spectrum of real-world scenarios, especially those that are less common but equally significant. Retrieval-augmented language models represent a promising approach for integrating rare or specialized knowledge into model outputs, effectively addressing challenges that cannot be resolved solely by scaling up training datasets (Kandpal et al., 2023).
@@
DGMs are also prone to adversarial vulnerability, often due to the presence of highly predictive but non-robust features that are used as shortcuts for prediction (Du et al., 2023a; Puli et al., 2023; Webson & Pavlick, 2022). This behavior poses a significant threat to various downstream scenarios, especially those of safety-critical applications (Poursaeed et al., 2021; Wang et al., 2023a). Several approaches to mitigate the effect of shortcut learning are based on model refinement or on dataset refinement, also known as data-centric approaches (Whang et al., 2021; Zha et al., 2025). In the former, work has been done towards improving robustness via adversarial training (Zou et al., 2023b), ensembling (Clark et al., 2019), contrastive learning (Choi et al., 2022), and the direct integration of prior knowledge (Ilyas et al., 2019). The latter includes improving the quality of the data used by large-scale models during training, such as through augmentation (Zhang et al., 2018), labeling (Kutlu et al., 2020)) and inference techniques—for example, employing prompt engineering (Wallace et al., 2019) or data slicing (Chung et al., 2019).
@@
However, in most applications, foundation models are often adapted to specific tasks and downstream datasets. Standard fine-tuning techniques often overemphasize the target task, leading to catastrophic forgetting (Thanh-Tung & Tran, 2020) and a loss in the general robustness of the upstream model (Suprem & Pu, 2022). Therefore, a significant challenge is to develop robust adaptation methods that adequately solve the target task but still maintain the beneficial robustness properties of the upstream model (e.g., robustness to distribution shifts of the target dataset) (Balaji et al., 2020; Du et al., 2023a; Han et al., 2021; Liu et al., 2020). These same issues come into play when developing smaller and more efficient models for the sake of economization of DGM inference and memory costs—which we discuss in greater detail in Section 3. In this context, it is important to develop robust distillation methods that do not sacrifice the robustness of the model (Du et al., 2023b; Zi et al., 2021). We argue that a promising approach to obtain robust and interpretable models is to explicitly encode human priors into the training process (Ilyas et al., 2019).
@@ 2.2 Accounting for Implicit Assumptions
Current generative models often make use of implicit assumptions and inductive biases, largely driven by convenience or convention. These practices are either rarely questioned or widely-embraced for convenience even when known to be wrong (Zhao et al., 2018). As one example, the algorithms used in machine learning often assume that data are drawn independently. In reality, data points are often correlated, such as in time-series data or through repeated measurements from the same individual. As another example, most generative models assume that latent distributions can be modeled on simple topological structures. However, latent distributions typically benefit from more expressive approaches (Stimper et al., 2022), suggesting the assumptions of their simplicity may be ill-founded.
We argue that convenience should not be the driving factor behind modeling assumptions. While the impact of model misspecifications on downstream applications in DGMs are not yet well understood, we have preliminary evidence suggesting their effects are undesirable: In traditional statistical analyses, such misspecifications are observed to have immense impacts (Cardon & Palmer, 2003); more recently, models that rely heavily on the training data distribution have been observed to exhibit bias and decreased performance if not properly corrected by meaningful modeling assumptions (Fortuin, 2022).
@@
Recent major breakthroughs in deep generative models (DGMs) have primarily been achieved in settings where models could be trained on internet-scale data (OpenAI, 2023; Rombach et al., 2022). However, many real-world applications, such as drug design (Vamathevan et al., 2019), material engineering (Wei et al., 2019), personalized medicine (MacEachern & Forkert, 2021), and protein biochemistry (Bonetta & Valentino, 2020), often have much smaller datasets due to the high cost of data generation. In these areas, domain experts often possess troves of detailed prior knowledge, which could potentially be used to enable more data-efficient learning in generative AI models. Indeed, it has been shown in the context of VAEs that incorporating domain prior knowledge can significantly improve model performance (Fortuin et al., 2020; Jazbec et al., 2021) and even unlock their use for tasks that were previously impossible (Fortuin et al., 2019; Manduchi et al., 2021; 2022).
@@
There are multiple routes via which prior knowledge can be encoded in generative AI systems (Dash et al., 2022). One straightforward way to incorporate domain knowledge is in Bayesian settings through the choice of prior distribution; such distributions can explicitly encode known properties of the target data. For example, an informed prior can reflect physiological constraints in medicine or chemical properties in materials science (Sam et al., 2024), taking a step towards ensuring that the learned model aligns with real-world principles. Beyond priors, domain knowledge can guide architectural design by suggesting specialized network components or hierarchical structures that reflect known relationships within the data (Andreas et al., 2016b; Shen et al., 2019; Bronstein et al., 2021) or can encourage models to process data in a more human-like manner for the sake of interpretability (Andreas et al., 2016a; McCoy et al., 2020; Vu et al., 2023; Lu et al., 2023). Finally, constraints embedded in either the model specification or the training algorithm can further ground generative models in real-world processes, leading to improved performance and trustworthiness (Raissi et al., 2019; Ren et al., 2020; Dash et al., 2021; Mohan et al., 2023). Each of these approaches can equip our models with helpful inductive biases that aid data-efficient learning.
@@
While designing future models with domain-informed inductive biases holds great promise, it may not be so straightforward in practice. For example, while VAEs are Bayesian models and, therefore, offer a natural paradigm for specifying a prior distribution over their latent space, many other DGMs lack such explicit mechanisms for encoding prior information. We consider diffusion models as a concrete example. At first, it might seem that the diffusion process’s Gaussian sampling distribution is comparable to the Gaussian latent prior in a VAE, suggesting a straightforward route for specifying priors for these models. However, this property of the diffusion process arises from the central limit theorem rather than from precise knowledge about the nature of the underlying data-generating distribution. Recent works have attempted to enhance the space of diffusion priors through auxiliary dimensions (Pandey & Mandt, 2023; Singhal et al., 2023). Unfortunately, neither of those approaches offers nearly the same flexibility of prior specification as the Bayesian priors in the latent space of VAEs, so further research into priors for diffusion models is sorely needed. In general, there are several caveats and hurdles that must be considered when designing future models with domain-informed inductive biases. First, by definition, biases constrain or push our models towards certain solutions. If the underlying bias does not capture every facet of the real-world process—an especially common concern in areas like biology, where core mechanisms remain poorly understood—it may inadvertently limit the model’s expressivity or lead to systematic errors. In other words, it may cause models to fail to learn important patterns that fall outside the imposed structure. Moreover, adding constraints often introduces computational challenges: physically or biologically inspired restrictions might be non-differentiable or otherwise difficult to incorporate into standard training pipelines, leading to more complex optimization procedures or increasing computational overhead. Thus, while biasing models with domain knowledge can significantly improve data efficiency and performance, careful consideration of both the correctness of those biases and the technical feasibility of their implementation is essential.
2.3 Causal Generative Models
Going beyond learning mere statistical correlations and understanding how underlying factors influence the generative process is the main objective of learning a causal structure of data (Pearl & Mackenzie, 2018). Such knowledge can be used to reason about hypothetical scenarios in the world, understand the effect of interventions, and perform counterfactuals (Pearl, 2019), thus facilitating informed decision-making. Although there have been attempts to develop methods for learning the optimal generative structure of deep latent variable models from data (He et al., 2019; Manduchi et al., 2023), current generative models often neglect the underlying causal dependencies in their generative processes, making them prone to shortcut learning and spurious associations (Gururangan et al., 2018; McCoy et al., 2019).
Causal generative models have the potential to offer distribution-shift robustness, fairness, and interpretability (Schölkopf et al., 2021; Wang & Jordan, 2021). They are either focused on causal representation learning, which discovers causally related latent variables, or controllable counterfactual generation, which, instead, focuses on learning a mapping between data and known causal variables. For a detailed review of the topic, we refer to (Komanduri et al., 2024). Current open challenges include but are not limited to, scalable and robust causal discovery from observational data (Reizinger et al., 2023; Zhou et al., 2022; Montagna et al., 2024), identifiability of deep generative models under weaker forms of supervision (Ahuja et al., 2023; Locatello et al., 2020; von Kügelgen et al., 2024), lack of benchmark datasets and metrics to evaluate counterfactual quality (Monteiro et al., 2023), strong assumptions that are often violated in real-world applications (Komanduri et al., 2024), and, finally, the integration of diffusion models, a field that is currently under-explored but has tremendous growth potential (Mittal et al., 2021; Pandey et al., 2022; Sanchez & Tsaftaris, 2022; Sanchez et al., 2023). We suggest that the integration of causal principles in DGMs could pave the way for the development of more robust, interpretable, and actionable generative AI systems (Zhou et al., 2023).
2.4 Foundation Models for Heterogeneous Data Types

### v4

@@ 1 Introduction
The past few years have demonstrated the immense potential of large-scale generative models to create powerful AI tools capable of impacting society profoundly. Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Rae et al., 2021) and their dialogue agents, such as ChatGPT (OpenAI, 2023) and Llama 3 (Grattafiori et al., 2024) have enabled the development of highly effective text generation systems that produce coherent, contextually relevant, and user-tailored outputs across a wide range of use cases. Similarly, advancements in diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) have led to groundbreaking advancements in image synthesis tasks, such as large-scale text-to-image generation (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Esser et al., 2024). These successes show that highly effective AI systems can be built using a relatively straightforward recipe: combining simple generative modeling paradigms (Larochelle & Murray, 2011; Sohl-Dickstein et al., 2015) with successful network architectures (Vaswani et al., 2017; Dosovitskiy et al., 2020; Ronneberger et al., 2015), training on large-scale datasets, and the incorporation of preferences via human feedback (Ouyang et al., 2022; Ziegler et al., 2019). The impact of generative AI has not been limited to text and image generation applications. It has fueled accelerated progress across a variety of research fields and practical applications, spanning from biology (Jumper et al., 2021) to weather forecasting (Ravuri et al., 2021), code generation (Chen et al., 2021b; Li et al., 2022b), video creation (Yang et al., 2023c; Ho et al., 2022b; Singer et al., 2022; Brooks et al., 2024), audio synthesis (Borsos et al., 2023; Liu et al., 2023a), and even artistic and musical composition (Huang et al., 2023b).
@@
With the current advancements and excitement surrounding generative AI, a question naturally arises: Are we on the brink of an AI utopia? Are we close to developing what we might call a perfect generative model? For the purpose of this survey, we define such a model as a single system that (i) can approximate the joint data distribution of any modality, (ii) provides calibrated uncertainty assessments and demonstrates causal consistency, and (iii) delivers controllable outputs that satisfy stringent requirements on robustness, safety, efficiency and societal alignment. We argue in this paper that the answer is a resounding no; rather, the realization of such a model, one that would fundamentally transform the field of AI is still hampered by substantial theoretical, practical, and ethical challenges, and incremental advances alone are unlikely to close the gap in the near term.
@@
Amidst the excitement and anticipation surrounding this new wave of Deep Generative Models (DGMs), [^1] In this paper, we take the term Generative AI to refer to systems whose core component is a large-scale DGM. For brevity, we refer to this class and its instances simply as DGMs. it is easy to overlook the new set of challenges they introduce. Unlike many of the traditional machine learning models, DGMs generate outputs in very high-dimensional spaces, which introduces several technical complexities. These include significantly increased computational demands, a need for larger datasets to accurately capture the underlying data distribution, and challenges in effectively evaluating and interpreting the generated outputs (Tong et al., 2024). And while significant progress has been made in improving interpretability and computational efficiency for traditional models (Marcinkevics & Vogt, 2020), these existing methods are frequently ill-suited for DGMs (Singh et al., 2024), at least in part because of the complex and high-dimensional nature of their outputs. Consequently, there is a pressing need for the development of a new set of techniques and tools tailored to these models, particularly to enable efficient inference, interpretability and quantization. These challenges lead us to conclude that scaling up current paradigms is not in isolation the ultimate path towards a perfect generative model. While increasing model size and training data can enhance performance on benchmarks (Hoffmann et al., 2022), it does little to address the fundamental shortcomings of DGMs, such as their inefficiency, lack of inclusivity, limited transparency, and barriers to usability—particularly in high-stakes domains where reliability and fairness are paramount.
@@
By presenting a broader roadmap of the current state and open challenges in generative AI, this paper offers an integrated entry point for researchers and practitioners. Although many existing surveys offer in-depth reviews of specific subfields, such as robustness, causal modeling, or modality-specific techniques, they are often narrowly focused, making it difficult to grasp the collective landscape. In contrast, we put forward high-level insights across domains, highlight emerging research directions, and guide readers to foundational and topic-specific work. In doing so, our goal is to foster the development of generative AI systems that are more robust, inclusive, and accessible.
This paper emerged as a result of the Dagstuhl Seminar on Challenges and Perspectives in Deep Generative Modeling [^2] https://www.dagstuhl.de/23072 held in Spring 2023.
@@ 2 Expanding Scope and Adaptability
Sub-challenge Typical failure mode Promising mitigation avenues Reference materials Research-question candidates
Robust generalisation to OOD inputs (§2.1) Large performance drop on unseen domains Retrieval-augmented generation (RAG); Group DRO training (e.g., Sagawa* et al., 2020) Datasets:
WILDS
; ImageNet-C
Surveys:
Shen et al. (2021); Yang et al. (2023b); Li et al. (2023c)
Q1: Can retrieval-augmented generators help close the WILDS gap without model retraining?
Q2: Can we provide an inductive bias for foundation models via architectural modifications or training objectives that predisposes them to accurately capture tail events?
Resilience to adversarial perturbations (§2.1) Imperceptible noise fools the model Gaussian noise smoothing; Certified defense methods; Adversarial finetuning for diffusion Benchmarks:
RobustBench Leaderboard
; NSFW Adversarial Benchmark
Surveys:
Sun et al. (2023b)
Q1: Will certifiably-robust diffusion sampling algorithms scale to large models, e.g., ones trained on ImageNet?
Q2: How can we unify adversarial training across different modalities for multi-modal models in a single framework?
Mitigating learning of spurious correlations (§2.1 / 2.2) Model predicts based on background cues, not capturing meaningful relationships Counterfactual data augmentation; Invariant Risk Minimization-based approaches Datasets:
Waterbirds
; Colored-MNIST
Surveys:
Ye et al. (2024)
Q1: How can we reliably detect and quantify hidden spurious cues encoded by foundation-model features?
Q2: Which causal probes best expose hidden biases in high-dimensional latent spaces?
Capturing causal dependencies (§2.2) Models generate statistically plausible but causally impossible outcomes SCM-guided loss functions; Interventional training objectives Benchmarks:
CausalBench
Surveys:
Komanduri et al. (2024)
Q1: How can causal invariance objectives be integrated into generative model training to improve robustness to distribution shifts?
Q2: Can tractable surrogate objectives approximate interventional likelihood, enabling scalable causal DGM training?
Accounting for implicit assumptions (§2.3) (Implicitly) assumed characteristics of data-generating distribution do not persist under domain shift Domain-expert knowledge integration; Statistical assumption testing (e.g., independence, stationarity checks) Q1: How can we quantify the degree of modeling-assumption violations in a trained generative model before deploying it?
Q2: What learning objectives are most robust when the true data-generating process lies outside the model family?
Cross-modal transfer in specialized domains (§2.4) Failure/inability to link signals across modalities (e.g. ECG $\leftrightarrow$ notes) Contrastive multimodal pre-training (e.g., Raghu et al., 2022) Datasets:
MMIST-CCRCC
; GMAI-MMBench
Surveys:
Shaik et al. (2024)
Q1: How can physiological constraints be incorporated into multi-modal pre-training objectives for medical domains?
Q2: Which training/fine-tuning methods can achieve sufficient cross-modal alignment in low-resource clinical settings?
Table 1: Research challenges summary table - Expanding Scope and Adaptability
@@ 2.1 Generalization and Robustness
To ensure reliability across various domains, DGMs must generalize effectively under shifts to the data-generating distribution of inputs, often referred to as out-of-distribution (OOD) robustness, and be resilient to minor variations in the input, a necessary component of the broader notion of adversarial robustness. Without proper generalization, generative models may produce unrealistic or biased outputs, [^3] Here, we use the terminology biased outputs to refer to systematic deviations in model outputs caused by imbalances or inaccuracies in the training data and/or modeling process. These outputs then do not accurately reflect the true underlying data distribution or are skewed in ways that perpetuate inaccuracies, stereotypes, or unfair conceptions about certain outcomes. limiting their practical utility and trustworthiness in real-world applications.
@@
While large-scale generative models show some promise in achieving OOD robustness (Wang et al., 2023a), these models still face challenges in accurately capturing rare events or responding to adversarial inputs (Zhu et al., 2024), a difficulty that lies in effectively modeling the tail of information (Kandpal et al., 2023), i.e., the information that appears rarely or only once in the dataset used to (pre)train the model. This limitation indicates a gap in their ability to fully represent the vast and diverse spectrum of real-world scenarios, especially those that are less common but equally significant. Retrieval-augmented language models represent a promising approach for integrating rare or specialized knowledge into model outputs, effectively addressing challenges that cannot be resolved solely by scaling up training datasets (Kandpal et al., 2023). In the vision domain, test-time approaches, such as Generalized Diffusion Adaptation (Tsai et al., 2024), present a promising avenue towards attaining OOD robustness.
@@
DGMs are also prone to adversarial vulnerability, often due to the presence of highly predictive but non-robust features that are used as shortcuts for prediction (Du et al., 2023a; Puli et al., 2023; Webson & Pavlick, 2022). This behavior poses a significant threat to various downstream scenarios, especially those of safety-critical applications (Poursaeed et al., 2021; Wang et al., 2023a). Several approaches to mitigate the effect of shortcut learning are based on model refinement or on dataset refinement, also known as data-centric approaches (Whang et al., 2021; Zha et al., 2025). In the former, work has been done towards improving robustness via adversarial training (Zou et al., 2023b; Choi et al., 2025), feature masking during training (Asgari et al., 2022), ensembling (Clark et al., 2019), contrastive learning (Choi et al., 2022), and the direct integration of prior knowledge (Ilyas et al., 2019). The latter includes improving the quality of the data used by large-scale models during training, such as through augmentation (Zhang et al., 2018), labeling (Kutlu et al., 2020)) and inference techniques—for example, employing prompt engineering (Wallace et al., 2019) or data slicing (Chung et al., 2019).
@@
However, in most applications, foundation models are often adapted to specific tasks and downstream datasets. Standard fine-tuning techniques often overemphasize the target task, leading to catastrophic forgetting (Thanh-Tung & Tran, 2020) and a loss in the general robustness of the upstream model (Suprem & Pu, 2022). Therefore, a significant challenge is to develop robust adaptation methods that adequately solve the target task but still maintain the beneficial robustness properties of the upstream model (e.g., robustness to distribution shifts of the target dataset) (Balaji et al., 2020; Du et al., 2023a; Han et al., 2021; Liu et al., 2020). These same issues come into play when developing smaller and more efficient models for the sake of economization of DGM inference and memory costs—which we discuss in greater detail in Section 3. In this context, it is important to develop robust distillation methods that do not sacrifice the robustness of the model (Du et al., 2023b; Zi et al., 2021). We argue that two particularly promising approaches to obtaining robust and interpretable models are embedding causal structure and explicitly encoding human priors into the training process—topics we examine in the following sections.
2.2 Causal Generative Models
Going beyond learning mere statistical correlations and understanding how underlying factors influence the generative process is the main objective of learning a causal structure of data (Pearl & Mackenzie, 2018). Structural Causal Models (SCMs) provide the mathematical foundation for this endeavor, representing causal relationships through directed acyclic graphs paired with structural equations that encode how variables causally influence one another. Such knowledge can be used to reason about hypothetical scenarios in the world, understand the effect of interventions, and perform counterfactuals (Pearl, 2019), thus facilitating informed decision-making. Although there have been attempts to develop methods for learning the optimal generative structure of deep latent variable models from data (He et al., 2019; Manduchi et al., 2023), current generative models often neglect the underlying causal dependencies in their generative processes, making them prone to shortcut learning and spurious associations (Gururangan et al., 2018; McCoy et al., 2019).
Causal generative models have the potential to offer distribution-shift robustness, fairness, and interpretability (Schölkopf et al., 2021; Wang & Jordan, 2021). They are either focused on causal representation learning, which discovers causally related latent variables, or controllable counterfactual generation, which, instead, focuses on learning a mapping between data and known causal variables. For a detailed review of the topic, we refer to (Komanduri et al., 2024). Current open challenges include but are not limited to, scalable and robust causal discovery from observational data (Reizinger et al., 2023; Zhou et al., 2022; Montagna et al., 2024), identifiability of DGMs under weaker forms of supervision (Ahuja et al., 2023; Locatello et al., 2020; von Kügelgen et al., 2024), lack of benchmark datasets and metrics to evaluate counterfactual quality (Monteiro et al., 2023), strong assumptions that are often violated in real-world applications (Komanduri et al., 2024), and, finally, the integration of diffusion models, a field that is currently under-explored but has tremendous growth potential (Mittal et al., 2021; Pandey et al., 2022; Sanchez & Tsaftaris, 2022; Sanchez et al., 2023). We suggest that the integration of causal principles in DGMs could pave the way for the development of more robust, interpretable, and actionable generative AI systems (Zhou et al., 2023).
@@ 2.3 Accounting for Implicit Assumptions
Current generative models often make use of implicit assumptions and inductive biases. Many of these, such as translational equivariance in CNNs or locality in audio diffusion models, are principled and empirically validated. Others, however, persist mainly for computational convenience [^4] By “convenience” we mean design choices adopted because they are easy to implement or tractable to optimize, not because they have been shown to match the true structure of the data. or historical precedent, even when their validity for specific applications remains unexamined or is blatantly known to be wrong (Zhao et al., 2018). As one example, the algorithms used in machine learning often assume that data are drawn independently. In reality, data points are often correlated, such as in time-series data or through repeated measurements from the same individual (Jiang & Nguyen, 2007; Kirchler et al., 2023). As another example, most generative models assume that latent distributions can be modeled on simple topological structures. However, latent distributions typically benefit from more expressive approaches (Stimper et al., 2022), suggesting the assumptions of their simplicity may be ill-founded.
We argue that convenience should not be the driving factor behind modeling assumptions. While the impact of model misspecifications on downstream applications in DGMs are not yet well understood, we have preliminary evidence suggesting their effects are undesirable: In traditional statistical analyses, such misspecifications are observed to have immense impacts (Cardon & Palmer, 2003); more recently, empirical studies have revealed systematic biases in DGMs that may stem from inadequate modeling assumptions (Zhao et al., 2018), and models that rely heavily on the training data distribution have been observed to exhibit bias and decreased performance if not properly corrected by meaningful modeling assumptions (Fortuin, 2022).
@@
Recent breakthroughs in DGMs have primarily been achieved in settings where models could be trained on internet-scale data (OpenAI, 2023; Rombach et al., 2022). However, many real-world applications, such as drug design (Vamathevan et al., 2019), material engineering (Wei et al., 2019), personalized medicine (MacEachern & Forkert, 2021), and protein biochemistry (Bonetta & Valentino, 2020), often have much smaller datasets due to the high cost of data generation. In these areas, domain experts often possess extensive prior knowledge, which could potentially be leveraged to enable more data-efficient learning in generative AI models. Indeed, it has been shown in the context of VAEs that incorporating domain prior knowledge can significantly improve model performance (Fortuin et al., 2020; Jazbec et al., 2021) and even unlock their use for tasks that were previously impossible (Fortuin et al., 2019; Manduchi et al., 2021; 2022).
@@
There are multiple routes via which prior knowledge can be encoded in generative AI systems (Dash et al., 2022). One straightforward way to incorporate domain knowledge is in Bayesian settings through the choice of prior distribution; such distributions can explicitly encode known properties of the target data. For example, an informed prior can reflect physiological constraints in medicine or chemical properties in materials science (Sam et al., 2024), taking a step towards ensuring that the learned model aligns with real-world principles. Similarly, recent work in diffusion models employs heavy-tailed priors to model extreme or rare events (Pandey et al., 2025). Therefore, we need generative models that can learn mappings between domain-specific priors and the observed data distributions more flexibly (Albergo et al., 2023; Lipman et al., 2023). Beyond priors, domain knowledge can guide architectural design by suggesting specialized network components or hierarchical structures that reflect known relationships within the data (Andreas et al., 2016b; Shen et al., 2019; Bronstein et al., 2021) or can encourage models to process data in a more human-like manner for the sake of interpretability (Andreas et al., 2016a; McCoy et al., 2020; Vu et al., 2023; Lu et al., 2023). Finally, constraints embedded in either the model specification or the training algorithm can further ground generative models in real-world processes, leading to improved performance and trustworthiness (Raissi et al., 2019; Ren et al., 2020; Dash et al., 2021; Mohan et al., 2023). Each of these approaches can equip our models with helpful inductive biases that aid data-efficient learning.
@@
While designing future models with domain-informed inductive biases holds great promise, it also presents several challenges that must be carefully considered (Battaglia et al., 2018; Bronstein et al., 2021). For example, while VAEs are Bayesian models and, therefore, offer a natural paradigm for specifying a prior distribution over their latent space, many other DGMs lack such explicit mechanisms for encoding prior information. We consider diffusion models as a concrete example. At first, it might seem that the diffusion process’s Gaussian sampling distribution is comparable to the Gaussian latent prior in a VAE, suggesting a straightforward route for specifying priors for these models. However, this property of the diffusion process arises from the central limit theorem rather than from precise knowledge about the nature of the underlying data-generating distribution. Recent works have attempted to enhance the space of diffusion priors through auxiliary dimensions (Pandey & Mandt, 2023; Singhal et al., 2023), and through alternative structured or learned priors (Trippe et al., 2023; Wu et al., 2024). Unfortunately, those approaches do not offer the same degree of flexibility of prior specification provided by the Bayesian priors in VAEs’ latent space, highlighting the need for continued research into priors for diffusion models. More broadly speaking, by definition, biases constrain or push our models towards certain solutions (Mitchell, 1980). If the underlying bias does not capture every facet of the real-world process—an especially common concern in areas like biology, where core mechanisms remain poorly understood—it may inadvertently limit the model’s expressivity or lead to systematic errors. That is, models may fail to learn important patterns that fall outside the imposed structure (Ghassemi et al., 2020). Adding constraints also often introduces computational challenges: physically or biologically inspired restrictions might be non-differentiable or otherwise difficult to incorporate into standard training pipelines, leading to more complex optimization procedures or increasing computational overhead, for example when enforcing PDE/ODE constraints or discrete combinatorial rules (Cranmer et al., 2020; Li et al., 2023e). Thus, while biasing models with domain knowledge can significantly improve data efficiency and performance, careful consideration of both the correctness of those biases and the technical feasibility of their implementation is essential.
2.4 Foundation Models for Domain-specific and Heterogeneous Modalities